OpenStudy (anonymous):

The digits of a 2-digit integer are K and V. If 10*K+V=16, which of the following equals 61? A) 10*V+K B) 10/V +K C) 10/K +V D) (K*V)/ 10 E)10/ (K*V)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

16=10*1+6 then K=1 and V=6 10V+K=10*6+1=61

OpenStudy (anonymous):

gotit thanx